OpenStudy (anonymous):

Josh works at $10.73 per hour. If he worked 47 hours this week, what is his gross pay? Remember that any time worked over 40 hours is paid at time and a half.

OpenStudy (anonymous):

So you first Take care of the 40 hours So you have 40x(10.73)=429.20 Because 40 hours are paid at the rate of 10.73 now add to 429.20 the earnings of the 7 hours worked overtime: + 7x[10.73+(10.73/2)]= 7x(10.73+5.365)=7x(16.095)=112.665. So the answer is 429.20+112.665=541.865 Or if you round it is 541.87. Gross means you don't take out taxes.